A sample of 8 items has an average fat content of 18.6 grams and a standard deviation of 2.4 grams. Assuming a normal distribution, construct a 99 percent confidence interval for μ.
Answer #1

Solution :

sample size = n = 8

Degrees of freedom = df = n - 1 = 7

t\alpha /2,df = 3.499

Margin of error = E = t\alpha/2,df * (s /\sqrtn)

= 3.499 * (2.4 / \sqrt 8)

Margin of error = E = 3.0

The 99% confidence interval estimate of the population mean is,

\bar x - E < \mu < \bar x + E

18.6 - 3.0 < \mu < 18.6+ 3.0

15.6 < \mu < 21.6

(15.6 , 21.6)